My daughter has math homework i cannot help with.
Geometry- Straight segments
She needs to find out what "z" is...
Y is the midpoint. y= -1,5
X= -4,3
What is z?
Formula= x-x/2 y-y/2

Answer by venugopalramana(3286)

My daughter has math homework i cannot help with.
Geometry- Straight segments
She needs to find out what "z" is...
Y is the midpoint. y= -1,5
X= -4,3
What is z?
Formula= x-x/2 y-y/2
IF X IS(X1,X2)....Z IS (Z1,Z2)....THEN THEIR MID POINT Y (Y1,Y2)IS GIVEN BY
Y1=(X1+Z1)/2...........Y2=(X2+Z2)/2
SO WE GET
Y1=-1=(-4+Z1)/2
-4+Z1=-2
Z1=-2+4=2
-----------
Y2=5=(3+Z2)/2
3+Z2=10
Z2=7
---HENCE POINT Z IS (2,7)